OpenStudy (anonymous):

i really dont understand this equation -34=6x-42-8x please help me

OpenStudy (anonymous):

combine like terms on left side -34 = -2*x-42 Is it simpler now?

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

solve for x in the equation -34=6x-42-8x -34 = -2*x-42 add like terms that is 6x and -8x make -2x on the right hand side +42=-2x +42 add 42 to both sides to get rid of the 42 and leave -2x by itself. 8=-2x -4=x now just divide both sides by -2 which is x=-4 :